Find the time it takes for $9,800 to double when invested at an annual interest rate of 10%, compounded continuously. Round your answer to 4 decimal places.
Answer by mathmate(429)

Question:
Find the time it takes for $9,800 to double when invested at an annual interest rate of 10%, compounded continuously. Round your answer to 4 decimal places.

Solution:
Future value for continuous interest rate
F=Pe^(rt)
or
F/P=2=e^(rt)
Take natural log
log(2)=rt
t=log(2)/r=log(2)/0.1=6.931 years.